OpenStudy (anonymous):

solve u/6 - u/10 =1/6

OpenStudy (anonymous):

you can multiply the whole equation by 30

OpenStudy (anonymous):

u= 5/2

OpenStudy (anonymous):

Multiply by 30 on both left hand side and right hand side of the equation and then try to solve 30(u/6 - u/10) =30(1/6)

OpenStudy (anonymous):